OpenStudy (anonymous):

1/2+1/8+?=1 i need help to solve this problem

OpenStudy (anonymous):

1/2+1/8+x=1 Solve for x. x= 1-1/8-1/2 x=3/8

OpenStudy (anonymous):

Dont forget to fan :$

OpenStudy (anonymous):

thank you for your help

OpenStudy (anonymous):

hey